# Exploring and Exploiting the Asymmetric Valley of Deep Neural Networks

### v3

@@ 1 Introduction
Perturbation analysis around the local minima of DNNs [7, 60], i.e., the shape of the valleys they reside in, is a very popular research topic. The concept of flat minima was originally proposed by [27], who defines the size of the connected region around the minima where the loss remains relatively unchanged as flatness. Subsequent studies debate whether the flat or sharp minima could reflect the generalization ability [38, 43, 35, 55, 16, 12, 41, 3, 36]. The previous works constrain the valley shape to be symmetric, while recent work points out that not all DNN valleys are flat or sharp, and there also exist asymmetric valleys [24], which has not been systematically studied as far as we know.
@@
Aside from empirical observations, theoretical insights for our interesting findings are provided. We first declare that adding sign-consistent noise to parameters may have a larger probability of keeping activating the neurons or keeping the overwhelming score in classification tasks. Then, we show that the trace of the Hessian matrix along the sign-consistent direction is smaller, implying a flatter region. The above findings inspire applications in the fields of model fusion [47, 65, 32, 2]. This paper first explains why model aggregation based on pre-trained models often leads to performance improvements, i.e., the success of model soups [65], and then proposes constraining the sign of DNN parameters in federated learning [52, 68] to facilitate aggregation.
@@ 2 Related Works
Exploring the Valley Shape of DNNs. The valley around a minima solution has originally been viewed as flat or sharp [27], and the large batch size training may lead to sharp minima with poor generalization [38]. However, [12] declares that two solutions that are scale-invariant in performance may lie in regions with significantly different flatness. [43] verifies the findings of [12] by applying varying weight decay to small and large batch training, leading to results contrary to [38]. With the filter-normalized plots, [43] again observes that sharpness correlates well with generalization error. Later work also shows that the sharpness calculation should be dependent on the parameter scale [41] or the dataset [3]. Whether valley width reflects the generalization is still inconclusive. The proposal of asymmetric valley [24] further throws this debate into limbo, as valleys around minima could be flat on one side but sharp on the other side, which makes it more difficult to define flatness. [36] studies the effectiveness of two flat minima optimizers by visualizing the linear interpolation curves between them, which also display asymmetric valleys. However, they do not explain the existence of asymmetry valley, while this paper thoroughly studies the causes and implications of the unexplained asymmetry phenomenon.
@@
Model Fusion. Directly averaging two independent models may encounter a barrier due to training randomness or permutation invariance [4, 22, 2, 49]. However, if the two models are generated from separate training of a common pre-trained model, the model fusion may perform better than individual models, i.e., the model soups [56, 65]. We will explain the success of model soups based on the findings in our paper. Popular federated learning algorithms also take the parameter averaging process to fuse the individual models updated on isolated clients [52, 50]. A huge challenge is the Non-Independent and Identical Data distributions of data islands (Non-I.I.D. data) [68, 28], which could make local models too diverged to merge. Multiple regularization methods are proposed to align parameters before model fusion [46, 1, 45, 37]. We propose an effective regularization method that focuses on the sign of parameters, which is inspired by our interesting findings.
@@ 3 Basic Notations and Preliminaries
Our major tool is plotting the 1D error curve of DNNs following the formula $\theta_{f}+\lambda\epsilon$ (Fig. 2 (A)). $\theta_{f}$ denotes the converged model, and $\epsilon$ denotes a noise vector sampled from a specific distribution.
@@ 6.1 Explaining the Success of Model Soups
We experimentally verify our guess on two datasets, i.e., training VGG16BN with completely random initialization on CIFAR10, and training ResNet18 with pre-trained initialization from PyTorch on Flowers [57]. The datasets are first uniformly split into two partitions, and then two models with corresponding initialization are separately trained or fine-tuned for 50 epochs. The checkpoints in the $\{1,2,3,5,10,20,30,50\}$-th epoch are stored. For checkpoints in the specific epoch, we denote the two models on the two partitions as $\theta_{A}$ and $\theta_{B}$, respectively. The interpolation accuracy of $(1-\lambda)\theta_{A}+\lambda\theta_{B}$ on the test set is plotted in Fig. 11. The interpolation curve of VGG16BN on CIFAR10 indeed encounters a significant barrier, especially when the epoch is larger, e.g., $E=50$. In contrast, the interpolation surpasses the individual models on Flowers, which is attributed to the pre-trained ResNet18. As an explanation, we calculate the sign consistency ratio between $\theta_{A}$ and $\theta_{I}$, $\theta_{B}$ and $\theta_{I}$, and $\theta_{A}$ and $\theta_{B}$, denoted as “SSR-IA”, “SSR-IB”, and “SSR-AB”, respectively. $\theta_{I}$ means the initialization model. We also plot the gap of model interpolation and individual models when $\lambda=0.5$, i.e., $\text{Acc}(0.5\theta_{A}+0.5\theta_{B})-0.5(\text{Acc}(\theta_{A})+\text{Acc}(\theta_{B}))$. The right of Fig. 11 clearly shows that the sign consistency ratio could almost perfectly reflect the tendency of the interpolation gap. Notably, the sign consistency ratio of models on Flowers is higher than $0.95$, which means that fine-tuning the pre-trained model does not change the parameter signs a lot, which facilitates the following parameter interpolation.
Figure 11: The models fine-tuned from a pre-trained model have a higher sign consistency ratio.
@@ 6.2 Regularizing the Sign Change in Federated Learning
$\mathcal{L}^{k}=\mathcal{L}_{\text{ce}}^{k}+\gamma\left(\text{sgp}(\theta_{t})\sigma(\theta_{t}^{k})+\text{sgp}(-\theta_{t})\sigma(-\theta_{t}^{k})\right),$ (1)
@@
CIFAR-10 10.0 81.96 81.84 82.44 80.45 81.92 82.61
1.0 80.48 80.42 80.12 79.78 80.30 80.36
0.5 77.53 78.12 76.77 78.02 77.78 78.59
CINIC-10 10.0 76.09 76.58 76.25 76.37 76.84 77.29
1.0 72.17 72.25 72.06 73.12 73.46 74.44
0.5 68.42 69.11 68.55 69.01 70.14 70.66
Experimental studies are verified on CIFAR10 [39] and CINIC10 [8] that are commonly utilized in previous works [67, 49]. Decentralizing the training data of these datasets by a Dirichlet distribution could simulate the Non-I.I.D. scenes as in real-world FL. The Dirichlet alpha $\alpha$ is utilized to control the Non-I.I.D. level, with a smaller $\alpha$ representing a more rigorous heterogeneity between clients’ data. We set $\alpha\in\{10.0,1.0,0.5\}$ respectively. The number of clients is $100$, and the total communication round is $200$. During each round, a random set of $10\%$ clients participate in FL and every client takes $5$ epochs update on their individual data. After all communication rounds, we evaluate the model performance on a global test set on the server (i.e., the original test set of corresponding datasets). We select our hyperparameter $\gamma$ from $\{0.001,0.01,0.1\}$ and report the best results. For $\alpha=10.0$, i.e., a relatively I.I.D. scene, a smaller $\gamma$ is better. In contrast, $\gamma=0.1$ or $\gamma=0.01$ will be more proper for $\alpha=0.5$. The performance comparison results are listed in Tab. 1. FedSign could surpass the compared methods, which shows the positive effects of regularizing the sign change in FL.

### v4

@@ 1 Introduction
Perturbation analysis around the local minima of DNNs [7, 61], i.e., the shape of the valleys they reside in, is a very popular research topic. The concept of flat minima was originally proposed by [27], who defines the size of the connected region around the minima where the loss remains relatively unchanged as flatness. Subsequent studies debate whether the flat or sharp minima could reflect the generalization ability [37, 42, 35, 56, 16, 12, 40, 3]. The previous works constrain the valley shape to be symmetric, while recent work points out that not all DNN valleys are flat or sharp, and there also exist asymmetric valleys [24], which has not been systematically studied as far as we know.
@@
Aside from empirical observations, theoretical insights for our interesting findings are provided. We first declare that adding sign-consistent noise to parameters may have a larger probability of keeping activating the neurons or keeping the overwhelming score in classification tasks. Then, we show that the trace of the Hessian matrix along the sign-consistent direction is smaller, implying a flatter region. The above findings inspire applications in the fields of model fusion [46, 66, 32, 2]. This paper first explains why model aggregation based on pre-trained models often leads to performance improvements, i.e., the success of model soups [66], and then proposes constraining the sign of DNN parameters in federated learning [53, 70, 49] to facilitate aggregation.
@@ 2 Related Works
Exploring the Valley Shape of DNNs. The valley around a minima solution has originally been viewed as flat or sharp [27], and the large batch size training may lead to sharp minima with poor generalization [37]. However, [12] declares that two solutions that are scale-invariant in performance may lie in regions with significantly different flatness. [42] verifies the findings of [12] by applying varying weight decay to small and large batch training, leading to results contrary to [37]. With the filter-normalized plots, [42] again observes that sharpness correlates well with generalization error. Later work also shows that the sharpness calculation should be dependent on the parameter scale [40] or the dataset [3]. Whether valley width reflects the generalization is still inconclusive. The proposal of asymmetric valley [24] further throws this debate into limbo, as valleys around minima could be flat on one side but sharp on the other side, which makes it more difficult to define flatness. This paper thoroughly studies the causes and implications of the unexplained asymmetry phenomenon.
@@
Model Fusion. Directly averaging two independent models may encounter a barrier due to training randomness or permutation invariance [4, 22, 2, 50]. However, if the two models are generated from separate training of a common pre-trained model, the model fusion may perform better than individual models, i.e., the model soups [57, 66]. A recent work [68] finds that resolving sign conflicts when merging multiple task-specific models is neccessary, which is most related to our current work. We will explain the success of model soups based on the relation between the asymmetric valley and the sign consistency of model parameters. Popular federated learning algorithms also take the parameter averaging process to fuse the individual models updated on isolated clients [53, 51]. A huge challenge is the Non-Independent and Identical Data distributions of data islands (Non-I.I.D. data) [70, 28], which could make local models too diverged to merge. Multiple regularization methods are proposed to align parameters before model fusion [45, 1, 44, 36]. We propose an effective regularization method that focuses on the sign of parameters, which is inspired by our interesting findings.
@@ 3 Basic Notations and Preliminaries
Our major tool is plotting the 1D error curve of DNNs following the formula $\theta_{f}+\lambda\epsilon$ (Fig. 2 (A)). $\theta_{f}$ denotes the converged model, and $\epsilon$ denotes a noise vector sampled from a specific distribution. More about the visualization of DNN loss landscape could be found in [42, 48].
@@ 6.1 Explaining the Success of Model Soups
We experimentally verify our guess on two datasets, i.e., training VGG16BN with completely random initialization on CIFAR10, and training ResNet18 with pre-trained initialization from PyTorch on Flowers [58]. The datasets are first uniformly split into two partitions, and then two models with corresponding initialization are separately trained or fine-tuned for 50 epochs. The checkpoints in the $\{1,2,3,5,10,20,30,50\}$-th epoch are stored. For checkpoints in the specific epoch, we denote the two models on the two partitions as $\theta_{A}$ and $\theta_{B}$, respectively. The interpolation accuracy of $(1-\lambda)\theta_{A}+\lambda\theta_{B}$ on the test set is plotted in Fig. 11. The interpolation curve of VGG16BN on CIFAR10 indeed encounters a significant barrier, especially when the epoch is larger, e.g., $E=50$. In contrast, the interpolation surpasses the individual models on Flowers, which is attributed to the pre-trained ResNet18. As an explanation, we calculate the sign consistency ratio between $\theta_{A}$ and $\theta_{I}$, $\theta_{B}$ and $\theta_{I}$, and $\theta_{A}$ and $\theta_{B}$, denoted as “SSR-IA”, “SSR-IB”, and “SSR-AB”, respectively. $\theta_{I}$ means the initialization model. We also plot the gap of model interpolation and individual models when $\lambda=0.5$, i.e., $\text{Acc}(0.5\theta_{A}+0.5\theta_{B})-0.5(\text{Acc}(\theta_{A})+\text{Acc}(\theta_{B}))$. The right of Fig. 11 clearly shows that the sign consistency ratio could almost perfectly reflect the tendency of the interpolation gap. Notably, the sign consistency ratio of models on Flowers is higher than $0.95$, which means that fine-tuning the pre-trained model does not change the parameter signs a lot, which facilitates the following parameter interpolation. For better illustration and fair comparison, we also replace the pre-trained ResNet18 with a random initialized ResNet18, and we find the plots shown in Fig. 11 tend to be like the results of the random initialized VGG16BN. The previous work [68] points out that disagreement on the sign of a given parameter’s values across models is a major source of interference during model merging. Although our finding is similar to the previous work, the motivation and specific explanation differs a lot.
Figure 11: The models fine-tuned from a pre-trained model have a higher sign consistency ratio. The top two sub-figures do not utilize pre-trained models, while the last one utilizes pre-trained ResNet18.
@@ 6.2 Regularizing the Sign Change in Federated Learning
$\mathcal{L}^{k}=\mathcal{L}_{\text{ce}}^{k}-\gamma\left(\text{sgp}(\theta_{t})\sigma(\theta_{t}^{k})+\text{sgp}(-\theta_{t})\sigma(-\theta_{t}^{k})\right),$ (1)
@@
CIFAR-10 10.0 81.53 $\pm$ 0.17 81.84 82.44 80.45 81.92 82.59 $\pm$ 0.09
1.0 80.54 $\pm$ 0.11 80.42 80.12 79.78 80.30 80.76 $\pm$ 0.14
0.5 77.69 $\pm$ 0.21 78.12 76.77 78.02 77.78 78.41 $\pm$ 0.35
CINIC-10 10.0 75.74 $\pm$ 0.24 76.58 76.25 76.37 76.84 77.05 $\pm$ 0.14
1.0 72.19 $\pm$ 0.15 72.25 72.06 73.12 73.46 74.59 $\pm$ 0.20
0.5 68.24 $\pm$ 0.32 69.11 68.55 69.01 70.14 70.63 $\pm$ 0.16
Experimental studies are verified on CIFAR10 [38] and CINIC10 [8] that are commonly utilized in previous works [69, 50]. Decentralizing the training data of these datasets by a Dirichlet distribution could simulate the Non-I.I.D. scenes as in real-world FL. The Dirichlet alpha $\alpha$ is utilized to control the Non-I.I.D. level, with a smaller $\alpha$ representing a more rigorous heterogeneity between clients’ data. We set $\alpha\in\{10.0,1.0,0.5\}$ respectively. The number of clients is $100$, and the total communication round is $200$. During each round, a random set of $10\%$ clients participate in FL and every client takes $5$ epochs update on their individual data. After all communication rounds, we evaluate the model performance on a global test set on the server (i.e., the original test set of corresponding datasets). We select our hyperparameter $\gamma$ from $\{0.001,0.01,0.1\}$ and report the best results. For $\alpha=10.0$, i.e., a relatively I.I.D. scene, a smaller $\gamma$ is better. In contrast, $\gamma=0.1$ or $\gamma=0.01$ will be more proper for $\alpha=0.5$. The performance comparison results are listed in Tab. 1. For FeaAvg and our proposed FedSign, we rerun the experimental studies five times and list the standard deviation of accuracies, showing that the accuracy doesn’t fluctuate very much. FedSign could surpass the compared methods, which shows the positive effects of regularizing the sign change in FL.
@@ 8 Conclusion
Acknowledgements
This work is partially supported by National Science and Technology Major Project (2022ZD0114805), NSFC (62376118, 62006112, 62250069, 61921006), Collaborative Innovation Center of Novel Software Technology and Industrialization. Professor De-Chuan Zhan is the corresponding author.
@@ 32 P = softmax(Z, axis=1)
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract and introduction clearly list the paper’s scope and contributions. The contributions are summarized in Sect. 1.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The limitations and future works are presented in Sect. 7.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate ”Limitations” section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: We do not provide strict theoretical proof because it is complex for DNNs. We only provide theoretical insights to explain the observations in our paper, which does not need too many assumptions and complete proof. The theoretical insights are provided in Sect. 5 and Appendix D.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We report our experimental details and the phenomena in our paper are easy to reproduce. The details are provided in Appendix A. A demo code to reproduce our finding is provided in Appendix D.3 and Code 2. A pseudo-code for our application to federated learning is also provided in Algo. 1.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We provide core codes and a demo code to reproduce the observed phenomena in our paper. We do not provide codes with external links. The demo code is in Code 1 and Code 2.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: Experimental settings and details are provided in the Appendix A.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [N/A]
  34. Justification: Our paper presents interesting phenomena about DNNs, which do not need error bars. However, we provide multiple groups of experimental observations to verify that the phenomena are common across various conditions. For example, multiple groups of experimental studies are provided to support our finding, e.g., the plots in Appendix B.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer ”Yes” if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [N/A]
  39. Justification: The experimental studies in our paper do not need too much computation budget, which could be reproduced on mainstream devices.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: Our work satisfies the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: Our work has no societal impact because we focus on studying the basic properties of DNNs.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: Our paper does not pose such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [N/A]
  59. Justification: The dataset and pre-trained used in the work are publicly available, and we were unable to find the license for the dataset we used.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: We do not release new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.